# GHOST: Hallucination-Inducing Image Generation for Multimodal LLMs

### v1

@@ Appendix F GHOST Additional Results
In this section, we provide extended qualitative and quantitative results to complement the main paper. Figures 15, 16, and 17 present qualitative examples of GHOST applied to Qwen2.5, LLava-1.6, and GLM-4.1Thinking, respectively, showing the original images alongside GHOST-generated samples and the corresponding model responses. Tables 16, 17, and 18 report detailed sample statistics for each object category across the three victim models, including the number of generated, filtered, and successful samples. Finally, Tables 20, 20, and 21 provide class-wise transferability results, highlighting how GHOST-induced hallucinations on one victim model transfer to others.
@@
Table 19: Class-wise transferability with Qwen2.5 as the victim model.
Object LLaMA LLava-1.6 GLM
@@
Table 20: Class-wise transferability with Llava-1.6 as the victim model.
Object LLaMA Qwen2.5 GLM
@@
Appendix G Finetuning
A growing body of work has shown that models robustness can be substantially improved through fine-tuning on synthetically generated or edited images (Augustin et al., 2025; Parast et al., 2025; Wang et al., 2024b; Wu et al., 2025; Azizmalayeri et al., 2022; Zhang et al., 2025b; Mirzaei et al., 2025; Bai et al., 2025b).
G.1 Fine-tuning protocol.
We fine-tune Qwen2.5VL-7B using images generated by our GHOST procedure (targeting Qwen2.5VL -7B) to assess whether such counterfactual samples improve model performance. We include this as a proof-of-concept demonstration, with scaling to larger datasets and all classes reserved for future work. We adopt LoRA (Hu et al., 2022) for parameter-efficient adaptation; the exact hyperparameters are listed in Table 22.
G.2 Dataset construction.
[…]
To increase label fidelity for the synthesized positives, we relied on the COCO annotations , and filtered candidate generations by ranking their CLIP embeddings against a compositional text embedding of the target object (Sec. E.2), keeping the highest-scoring images to give inputs to SD unCLIP model. This process improves the likelihood that the positive set truly depicts the intended object while keeping the data domain consistent (synthetic vs. synthetic) across classes. Examples of positive samples generated are shown in Figure 18.
G.3 Model selection
We use POPE on the COCO training split as our checkpointing criterion. After each epoch, we compute the POPE Random setting on 1,000 samples and retain the checkpoint with the highest F1 score.
G.4 Model Evaluation
We evaluate the fine-tuned model on our generated images from other models (victim models) and on POPE using the COCO validation set. Specifically, we use the POPE Adversarial, Random, and Popular settings with maximum samples of $n=3000$ samples. The results are available in the main paper. To evaluate how much fine-tuning impacted the overall performance of the MLLM, we evaluated the model on VQA and captioning tasks. For the VQA task, we leverage the small VQAv2 dataset from the Hugging Face implementation [^1] https://huggingface.co/datasets/merve/vqav2-small. The dataset consists of an image–question pair and a single-word ground-truth answer. We compute accuracy for both the base model and the fine-tuned model on this task to assess the impact of fine-tuning. For the captioning task, we leverage COCO validation images and captions as ground truth. We generate captions for all COCO validation images with the base model and the fine-tuned model using the prompt: “Write a short caption for the given image.” Then we compute BERTScore (Zhang* et al., 2020) using the generated captions and COCO annotations as ground truth.
Table 22: Fine-tuning hyperparameters and LoRA configuration.
LoRA configuration
Rank ($r$) 8
Alpha 32
Dropout 0.05
Training setup
Optimizer Adam
Learning rate $5\times 10^{-6}$
Epochs 15
Batch size 16
Warmup ratio 0.10
Scheduler Cosine with warmup

### v2

@@ Appendix I Object Detector Usage
In this section, we provide extended qualitative and quantitative results to complement the main paper. Figures 18, 19, and 20 present qualitative examples of GHOST applied to Qwen2.5, LLava-1.6, and GLM-4.1Thinking, respectively, showing the original images alongside GHOST-generated samples and the corresponding model responses. Tables 24, 25, and 26 report detailed sample statistics for each object category across the three victim models, including the number of generated, filtered, and successful samples. Finally, Tables 27, 28, and 29 provide class-wise transferability results, highlighting how GHOST-induced hallucinations on one victim model transfer to others.
@@
Table 27: Class-wise transferability results with Qwen2.5 as the victim model.
Object LLaMA 3.2 LLava-1.6 GLM-4.1
@@
Table 28: Class-wise transferability results with Llava-1.6 as the victim model.
Object LLaMA 3.2 Qwen2.5 GLM-4.1